# On the Viability of Open-Source Financial Rails: Economic Security of Permissionless Consensus

### v1

@@ 4.3 Recovery
The following defines a recovery oracle sufficient for Stubborn Nakamoto’s recovery. The recovery oracle’s capabilities should be limited to those the community can plausibly implement. While we do not formally model the community’s capabilities to achieve consensus, we bear in mind that the community must have been able to initiate the protocol and provide the genesis block (potentially through a costly process) but is unlikely to be able to maintain up-to-date information on all local views of all honest nodes. Therefore, our definition of the recovery oracle requires only the knowledge of the view of a single honest node.
@@
A call to a recovery oracle at time $t$ updates the protocol’s setup by creating a new genesis block ${\sf B}_{G}$, which extends all first-confirmed blocks in the view of some arbitrarily selected honest node. The call also awakes honest nodes that halted to resume being active.
In other words, a call to the recovery oracle restarts the protocol in a similar manner to the initial launch of the Nakamoto protocol: the call announces a genesis block that any future chain must extend. All honest nodes ignore blocks that do not extend the new genesis block. The recovery oracle requires some contemporary knowledge of the state of honest nodes running the protocol, as it needs to maintain consistency with all finalized blocks. However, the recovery oracle is only required to know the state of one (arbitrarily selected) honest node and can thus be plausibly implemented by the community. [^55] The Stubborn Nakamoto protocol is specifically designed to allow a recovery oracle that uses only limited capabilities (that the community can plausibly implement); see 1 and the discussion below.
@@
An attacker with sufficient resources can cause the Stubborn Nakamoto protocol to lose liveness. Even after a call to the recovery oracle, the attacker can repeat its attack to cause the protocol to lose liveness again. Given the impossibility results mentioned above, such a loss of liveness is unavoidable for any protocol that maintains consistency. However, it is costly for the attacker to prevent liveness continuously: each liveness violation requires the attacker to create an attack chain that confirms a conflicting block, and all blocks in the attack chain are discarded after the call to the recovery oracle. Theorem 3 guarantees that when the attacker stops and Stubborn Nakamoto regains an honest majority, Stubborn Nakamoto recovers liveness (while maintaining consistency throughout).
The following lemma provides the key argument for consistency of Stubborn Nakamoto after recovery. A genesis block establishes agreement on an initial ledger. 1 implies that the recovery oracle establishes a ledger that maintains consistency with any transactions finalized by honest nodes before recovery.
@@
If some honest node $i$ finalizes block ${\sf B}$, then block ${\sf B}$ is the first-confirmed block in the view of all honest nodes. If some honest node $i$ finalizes block ${\sf B}$ and does not halt, then all honest nodes finalize block ${\sf B}$.
@@
Suppose that honest node $i$ confirmed block ${\sf B}$ at time $t$ and later finalized block ${\sf B}$. Consider honest node $j$ that never confirms a block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$. Then node $j$ must receive and confirm block ${\sf B}$ by time $t+\Delta$, and block ${\sf B}$ is the first-confirmed block in the view of $j$.
Consider honest node $j$ that at time $\tau$ confirms a block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$. Since $i$ did not confirm any block that conflicts with ${\sf B}$ by time $t+2\Delta$, we must have that $\tau>t+\Delta$. Because node $i$ confirms block ${\sf B}$ at time $t$, it must be that node $j$ receives and confirms block ${\sf B}$ by time $t+\Delta$. We have that $\tau>t+\Delta$, and therefore block ${\sf B}$ is the first-confirmed block in the view of $j$.
To prove the second part, suppose that honest node $i$ confirms block ${\sf B}$ at time $t$ and later finalizes block ${\sf B}$ and does not halt. Consider honest node $j$ that at time $\tau$ confirms a block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$. Because $i$ does not confirm any block that conflicts with ${\sf B}$ by time $t+4\Delta$, we must have that $\tau>t+3\Delta$. Again, because node $i$ confirms block ${\sf B}$ at time $t$, it must be that node $j$ receives and confirms block ${\sf B}$ by time $t+\Delta$. We have that $j$ confirms block ${\sf B}$ and waits for $2\Delta$ without confirming another conflicting block, and thus $j$ finalizes block ${\sf B}$. ∎
@@
By Theorem 2, the protocol satisfies liveness after it recovers. By Theorem 1, if nodes $i,j$ finalize blocks ${\sf B},{\sf B}^{\prime}$ of the same height before recovery, then ${\sf B}={\sf B}^{\prime}$. Likewise, by Theorem 1, if nodes $i,j$ finalize blocks ${\sf B},{\sf B}^{\prime}$ of the same height after recovery, then ${\sf B}={\sf B}^{\prime}$. Suppose that node $i$ finalizes block ${\sf B}$ before recovery. By 1, every honest node first-confirmed block ${\sf B}$. Therefore, after recovery, the block ${\sf B}$ is part of the chain of the new genesis block ${\sf B}_{G}$ and is finalized by any honest node $j$. ∎
@@ 4.4 Necessity of the recovery oracle
The recovery oracle serves several functions. It resolves some potential ambiguity among online nodes. [^56] Under corrupt majority, it is possible that some nodes finalize block ${\sf B}$ and other nodes halt without finalizing ${\sf B}$, and it is possible that some nodes first-confirm block ${\sf B}$ and other nodes first-confirm block ${\sf B}^{\prime}\neq{\sf B}$ (and no node finalizes either block ${\sf B}$ or block ${\sf B}^{\prime}$). It resolves potential ambiguity for late-joining nodes. It discards previous conflicting blocks by providing a new genesis block. Last, it informs nodes that halted that they should resume being active.
@@
As an illustration, consider a late-joining node $i$ that attempts to start mining Bitcoin. Several similar blockchains use Nakamoto consensus (for example, Dogecoin and Litecoin). Node $i$ can determine which of the chains is the Bitcoin blockchain by using the genesis block provided to it from the protocol’s setup.
However, there are also different blockchains that share Bitcoin’s genesis block and its blockchain up to a certain block. For example, Bitcoin Cash was created by a so-called “hard fork” that split off from Bitcoin and created a new blockchain. The initial portions of the ledgers of Bitcoin and Bitcoin Cash are identical, but they are independent systems, and the later parts of their ledgers diverge. Node $i$ needs more recent information to determine whether it connects to Bitcoin or Bitcoin Cash, as these are separate blockchains that share the same initial genesis block. [^57] The hard fork that created Bitcoin Cash included technical changes to the protocol. Bitcoin itself also adopted some technical changes at a later time. We assume that node $i$ arrives after both protocols have made these technical changes, and node $i$ must rely on an external resource to identify the correct chain. In particular, the total hash rate of each chain cannot provide the required identification, as node $i$ may be interested in connecting to the chain with a lower hash rate. The recovery oracle allows us to capture the fact that, in practice, users and nodes have access to external resources that can identify the two chains.
@@
Even under synchrony, there is no permissionless protocol88todo: 8needs to be defined in Section 2 or 3 that allows late joiners, always satisfies consistency, and for some $\rho>0$ and $\tau\geq 0$ satisfies liveness after time $T+\tau$ if from time $T$ onwards at least $1-\rho$ of the nodes are honest.
The intuition for the proof is that under any permissionless protocol, the attacker can create a parallel copy of the ledger. We show that an attacker can simulate the execution to the point of recovery and bring a late-joining node to adopt a parallel copy that creates a consistency violation. The attacks on Bitcoin discussed in Section 3 can be seen as instances of this general attack.
@@
5 Observations on Partial Synchrony and Proof-of-Stake
@@ A.2 Definitions
Because the Nakamoto consensus protocol requires randomness, a small probability of failure may be inevitable. The formal definition of liveness for the Nakamoto and Stubborn Nakamoto protocols bounds the failure probability by a security parameter $\lambda$. The protocol can tune its security parameter $\lambda$ by selecting a confirmation depth parameter $k$, with $k=\omega(\log\lambda)$.
For the sake of simplicity, we assume that an honest node that mines a block can include all the outstanding valid transactions (not included in the longest chain) given its local view. Let $T_{\rm conf}=T_{\rm conf}(\lambda,\Delta,n,\rho)$ be a function of the security parameter $\lambda$, network delay $\Delta$, number of nodes $n$, and fraction of adversary nodes $\rho$. We call $T_{\rm conf}$ the confirmation time.99todo: 9poly-length execution
@@
The protocol satisfies $T_{\rm conf}$-Liveness if there is some negligible666666 We say that ${\sf negl}:\mathbb{R}\to\mathbb{R}_{+}$ is a negligible function iff for any fixed polynomial function $p(\lambda)$, there exists $\lambda_{0}$ such that for any $\lambda>\lambda_{0}$, ${\sf negl}(\lambda)<1/p(\lambda)$. In other words, a negligible function is one that vanishes faster than any inverse-polynomial function. function ${\sf negl}(\cdot)$ such that with $1-{\sf negl}(\lambda)$ probability when some honest node receives a transaction ${\sf tx}$ at some time $t$, then by time $t+T_{\rm conf}$ all honest nodes’ finalized ledgers include ${\sf tx}$.
@@
Let $\nu=2pn\Delta$. The Network-Aware Honest Majority assumption is satisfied if $\nu<1/2$ and there exists some (arbitrarily small) constant $\phi\in(0,1)$ such that $(1-\rho)(1-\nu)\geq(1+\phi)\rho$.
The term $(1-\rho)(1-\nu)$ can be viewed as the fraction of honest mining power discounted by the network delay $\Delta$. Therefore, this assumption requires that the honest mining power, even after accounting for network delay, exceed corrupt mining power by a small constant margin.
@@ A.4 Proof of Theorem 2
Given an execution of the Stubborn Nakamoto protocol, we can construct an execution of the (underlying) Nakamoto [^68] The Nakamoto protocol uses the same confirmation depth $k$ as our Stubborn Nakamoto. protocol. If the two executions are identical, in the sense that any honest node maintains the same finalized ledger, then Stubborn Nakamoto’s properties follow from those of Nakamoto. We show that under the network-aware honest majority assumption the two executions are identical except for a negligible probability, and therefore the $T_{\rm conf}$-Liveness of Stubborn Nakamoto follows from the the $T_{\rm conf}$-Liveness and consistency of Nakamoto (Pass et al.,, 2017; Shi,, 2022).
First, we claim that the two executions cannot differ, except for a negligible probability, because Stubborn Nakamoto ignores any block that conflicts with a finalized block (or any block extending a conflicting block). In Nakamoto an honest node only needs to know the longest chains in its records: it always finalizes a prefix of some longest chain and mines to extend some longest chain. Consistency of Nakamoto guarantees that, except with negligible probability, if block ${\sf B}^{*}$ is finalized in some honest node’s view, then any chain that conflicts with ${\sf B}^{*}$ will never become the longest chain in any honest node’s view. In other words, if honest nodes ignore blocks that conflict with ${\sf B}^{*}$, it will not affect the node’s choice of blocks to mine or its finalized ledger.
The other reason the executions may differ is that an honest node $i$ may halt under Stubborn Nakamoto after some block ${\sf B}$ becomes at least $k$ blocks deep in some longest chain according to its local record at time $t$, and a conflicting block ${\sf B}^{\prime}\neq{\sf B}$ of the same height as ${\sf B}$ also becomes at least $k$ blocks deep is some longest chain according to its local record at some time $\tau\leq t+4\Delta$. We show that the probability of such event is negligible in $\lambda$.
@@
Let ${\sf chain}$ be the chain in node $i$’s local record in round $t$ containing ${\sf B}$ at least $k$-deep, and let ${\sf chain}^{\prime}$ be the chain in node $i$’s local record in round $t+4\Delta$ containing ${\sf B}^{\prime}$ at least $k$-deep. Let ${\sf B}_{\rm common}$ the last block shared in common by ${\sf chain}$ and ${\sf chain}^{\prime}$. Suppose that ${\sf B}_{\rm common}$ is mined in round $s-1$. It holds that all blocks in ${\sf chain}$ and ${\sf chain}^{\prime}$ after ${\sf B}_{\rm common}$ are mined after in round $s$ or later. Let $\tau=t-s$. Because the chains ${\sf chain}$ and ${\sf chain}^{\prime}$ differ by at least $k$ blocks, by total block upper bound (see Lemma A.3), it must be that $\tau>\frac{k}{2pn}$.
@@
Observe that it must be that ${\bf C}[s:t-\Delta]\leq{\bf A}[s:t+4\Delta]$. Otherwise, there would have been a block ${\sf B}$ mined by an honest node during a convergence opportunity between $[s,t-\Delta]$, and that block ${\sf B}$ would have to appear in both ${\sf chain}$ and ${\sf chain}^{\prime}$ (see Section 17.3 of Shi, 2022). Therefore, it suffices to prove that except with negligible in $\lambda$ probability, it must be that ${\bf C}[s:t-\Delta]>{\bf A}[s:t+4\Delta]$.
@@
${\bf C}[s:r-\Delta]>(1-\varepsilon)(1-2pn\Delta)\alpha(\tau-\Delta)\;,$
where $\alpha:=p\cdot(1-\rho)n$ denotes the expected number of honest nodes that mine a block in a given round. By adversarial block upper bound, (see Lemma A.2) we have that
${\bf A}[s:r+4\Delta]<(1+\epsilon^{\prime})\beta(\tau+4\Delta)\;,$
@@
${\bf C}[s:r-\Delta]$ $>(1-\varepsilon)(1-\nu)\alpha(\tau-\Delta)$
@@
$>(1+\varepsilon^{\prime})\beta(\tau+4\Delta)\;,$
@@
This implies that except with negligible in $\lambda$ probability we have that ${\bf C}[s:r-\Delta]>{\bf A}[s:r+4\Delta]$, completing the proof.

### v2

@@ 4.4 Recovery
The following defines a recovery oracle that enables the recovery of the Stubborn Nakamoto protocol. This definition formalizes the exact function the community (or another external source) must provide to the protocol. The recovery oracle is practical only if its capabilities are limited to those the community can plausibly implement. While we do not formally model the community’s capabilities to achieve consensus, we bear in mind that the community must have been able to initiate the protocol and provide the genesis block (potentially through a costly process) but is unlikely to maintain real-time knowledge of every honest node’s local view. Therefore, our definition of the recovery oracle relies solely on having access to the view of a single honest node.
@@
A call to a recovery oracle at time $t$ updates the protocol’s setup by creating a new genesis block ${\sf B}_{G}$, which extends all clear-certified blocks in the view of some arbitrarily selected honest node. The call also awakes honest nodes that halted.
In other words, a call to the recovery oracle restarts the protocol in a similar manner to the initial launch of the Nakamoto protocol: the call announces a genesis block that any future chain must extend. All honest nodes ignore blocks that do not extend the new genesis block. The recovery oracle requires some contemporary knowledge of the state of honest nodes running the protocol, as it needs to maintain consistency with all finalized blocks. However, the recovery oracle is only required to know the state of a single (arbitrarily selected) honest node. Thus, we argue that this recovery oracle can be plausibly implemented by the community. [^62] The Stubborn Nakamoto protocol is specifically designed to allow a recovery oracle that uses only limited capabilities (that the community can plausibly implement); see 2 and the discussion below.
@@
An attacker with sufficient resources can cause the Stubborn Nakamoto protocol to lose liveness. Even after a call to the recovery oracle, the attacker can repeat its attack to cause the protocol to lose liveness again. Given the impossibility results mentioned above, such a loss of liveness is unavoidable for any protocol that maintains consistency. However, it is costly for the attacker to continuously violate liveness: each liveness violation requires the attacker to create an attack chain that certifies or potentially certifies a conflicting block, and all blocks in the attack chain are discarded after the call to the recovery oracle. Theorem 5 guarantees that when the attacker stops and Stubborn Nakamoto regains an honest majority, Stubborn Nakamoto recovers liveness (while maintaining consistency throughout).
Remark 3.
Theorem 5 holds under the protocol modification described in Remark 1. If online nodes detect an attack chain and ignore it without halting—while invoking the recovery oracle—then late joiners will be directed to the correct ledger by the recovery oracle, ensuring consistency.
The following lemma provides the key argument for the consistency of Stubborn Nakamoto after recovery. A genesis block establishes agreement on an initial ledger. 2 implies that the recovery oracle establishes a ledger that maintains consistency with any transactions finalized by honest nodes before recovery.
@@
If some honest node $i$ finalizes block ${\sf B}$, then every honest node $j$ clear-certifies block ${\sf B}$. If some honest node $i$ finalizes block ${\sf B}$ and does not halt, then all honest nodes finalize block ${\sf B}$.
@@
Suppose that honest node $i$ certified block ${\sf B}$ at time $t_{1}$ and finalized block ${\sf B}$ at time $t_{2}=t_{1}+2\Delta$. Let $\sf Chain(\bar{{\sf B}})$ be a logest chain in the view of node $i$ at time $t$ such that ${\sf B}$ is at least $k$-deep in $\sf Chain(\bar{{\sf B}})$.
Suppose for the sake of contradiction that some honest node $j$ does not clear-certify ${\sf B}$. Node $j$ receives $\sf Chain(\bar{{\sf B}})$ by time $t+\Delta$. If $\sf Chain(\bar{{\sf B}})$ (or an extension of it) is a longest chain in the view of node $j$, then it must be that there is some block ${\sf B}^{\prime}$ such that ${\sf B}^{\prime}\not\sim{\sf B}$ and ${\sf B}^{\prime}\in\mathcal{C}_{j}^{t+\Delta}$. But then we have by 1 that ${\sf B}^{\prime}\in\mathcal{PC}_{i}^{t+2\Delta}(2\Delta)$, which is a contradiction to $i$ finalizing ${\sf B}$.
Suppose then that there is a longer chain $\sf Chain(\bar{{\sf B}}^{\prime})$ in the view of node $j$ at time $t+\Delta$. Suppose that block ${\sf B}$ does not appear $k$-deep in $\sf Chain(\bar{{\sf B}}^{\prime})$ (otherwise, we reach a contradiction using the previous argument). Then there must be a block ${\sf B}^{\prime}$ that is at least $k$-deep in $\sf Chain(\bar{{\sf B}}^{\prime})$ and ${\sf B}^{\prime}\not\sim{\sf B}$. Again, we have ${\sf B}^{\prime}\in\mathcal{C}_{j}^{t+\Delta}$, leading to a contradiction.
For the second part of the proof, suppose that honest node $i$ certified block ${\sf B}$ at time $t_{1}$, finalized block ${\sf B}$ at time $t_{2}=t_{1}+2\Delta$, and did not halt at time $t_{3}=t_{1}+4\Delta$. By the previous part of the proof, every honest node $j$ clear-certified block ${\sf B}$. Suppose for the sake of contradiction that some node $j$ did not finalize block ${\sf B}$. Observe that node $j$ certified block ${\sf B}$ by time $t_{1}+\Delta$, and therefore it must be that node $j$ halts without finalizing ${\sf B}$ at some time $t^{\prime}\leq t_{1}+3\Delta$. But this implies that there exists some ${\sf B}^{\prime}\in\mathcal{PC}_{j}^{t^{\prime}}(2\Delta)$ such that ${\sf B}^{\prime}\not\sim{\sf B}$, and by 1 we have that ${\sf B}^{\prime}\in\mathcal{PC}_{i}^{t^{\prime}+\Delta}(4\Delta)\subseteq\mathcal{PC}_{i}^{t_{3}}(4\Delta)$ in contradiction to $i$ not halting at time $t_{3}$.
@@
By Theorem 2, the protocol satisfies liveness after it recovers. By Footnote 56, if nodes $i,j$ finalize blocks ${\sf B},{\sf B}^{\prime}$ of the same height before recovery, then ${\sf B}={\sf B}^{\prime}$. Likewise, by Footnote 56, if nodes $i,j$ finalize blocks ${\sf B},{\sf B}^{\prime}$ of the same height after recovery, then ${\sf B}={\sf B}^{\prime}$. Suppose that node $i$ finalizes block ${\sf B}$ before recovery. By 2, every honest node clear-certified block ${\sf B}$. Therefore, after recovery, the block ${\sf B}$ is part of the chain of the new genesis block ${\sf B}_{G}$ and is finalized by any honest node $j$. ∎
@@ 4.5 Necessity of the recovery oracle
The recovery oracle serves several functions. It resolves some potential ambiguity among online nodes. [^63] Under corrupt majority, it is possible that some nodes finalize block ${\sf B}$ and other nodes halt without finalizing ${\sf B}$, and it is possible that some honest nodes certified block ${\sf B}$ and other honest nodes certified block ${\sf B}^{\prime}\not\sim{\sf B}$ (and no node finalizes either block ${\sf B}$ or block ${\sf B}^{\prime}$). It resolves potential ambiguity for late-joining nodes. It discards previous conflicting blocks by providing a new genesis block. Last, it informs nodes that halted that they should resume being active.
@@
As an illustration, consider a late-joining node $i$ that attempts to join Bitcoin and become a miner. Several similar blockchains use Nakamoto consensus (for example, Dogecoin and Litecoin), and node $i$ must determine which of the chains is the Bitcoin blockchain. Formally, nodes are able to identify the Bitcoin blockchain by using the genesis block provided to it from the protocol’s setup.
However, there are also different blockchains that share Bitcoin’s genesis block and its blockchain up to a certain block. For example, Bitcoin Cash was created by a so-called “hard fork” that split off from Bitcoin and created a new blockchain. The initial portions of the ledgers of Bitcoin and Bitcoin Cash are identical, but they are independent systems, and the later parts of their ledgers diverge. Node $i$ needs more recent information to determine whether it connects to Bitcoin or Bitcoin Cash, as these are separate blockchains that share the same initial genesis block. [^64] The hard fork that created Bitcoin Cash included technical changes to the protocol. Bitcoin itself also adopted some technical changes at a later time. We assume that node $i$ arrives after both protocols have made these technical changes, and node $i$ must rely on an external resource to identify the correct chain. In particular, the total hash rate of each chain cannot provide the required identification, as node $i$ may be interested in connecting to the chain with a lower hash rate. The recovery oracle allows us to capture the fact that, in practice, users and nodes are able to access resources (external to the protocol) that can identify the two chains.
@@
Even under synchrony, there is no permissionless protocol66todo: 6needs to be defined in Section 2 or 3 that allows late joiners, always satisfies consistency, and for some $\rho>0$ and $\tau\geq 0$ satisfies liveness after time $T+\tau$ if from time $T$ onwards at least $1-\rho$ of the nodes are honest.
The intuition for the proof is that under any permissionless protocol, the attacker can create a parallel copy of the ledger. We show that an attacker can simulate the execution to the point of recovery and bring a late-joining node to adopt a parallel copy that creates a consistency violation. Conceptually, such an attack is similar to a phishing attack which misdirects users to a malicious copy of a website. The attacks on Bitcoin discussed in Section 3 can be seen as instances of this general attack.
@@
5 Partial Synchrony and Proof-of-Stake
@@ A.2 Definitions
Because the Nakamoto consensus protocol requires randomness, a small probability of failure may be inevitable. The formal definition of liveness for the Nakamoto and Stubborn Nakamoto protocols bounds the failure probability by a security parameter $\lambda$. The protocol can tune its security parameter $\lambda$ by selecting a required block depth $k$, such that $k=\omega(\log\lambda)$.
For the sake of simplicity, we assume that an honest node that mines a block can include all the outstanding valid transactions (not included in the longest chain) given its local view. Let $T_{\rm conf}=T_{\rm conf}(\lambda,\Delta,n,\rho)$ be a function of the security parameter $\lambda$, network delay $\Delta$, number of nodes $n$, and fraction of adversary nodes $\rho$. We call $T_{\rm conf}$ the confirmation time.77todo: 7poly-length execution
@@
The protocol satisfies $T_{\rm conf}$-Liveness if there is some negligible737373 We say that ${\sf negl}:\mathbb{R}\to\mathbb{R}_{+}$ is a negligible function iff for any fixed polynomial function $p(\lambda)$, there exists $\lambda_{0}$ such that for any $\lambda>\lambda_{0}$, ${\sf negl}(\lambda)<1/p(\lambda)$. In other words, a negligible function is one that vanishes faster than any inverse-polynomial function. function ${\sf negl}(\cdot)$ such that with $1-{\sf negl}(\lambda)$ probability when some honest node receives a transaction ${\sf tx}$ at some time $t$, then by time $t+T_{\rm conf}$ all honest nodes’ finalized ledgers include ${\sf tx}$.
@@
Consider a polynomial length execution of Stubborn Nakamoto with a required block depth $k$ that is super-logarithmic function in the desired security parameter $\lambda$. Let $\nu=2pn\Delta$. The Network-Aware Honest Majority assumption is satisfied if $\nu<1/2$ and there exists some (arbitrarily small) constant $\phi\in(0,1)$ such that
$(1-\rho)(1-\nu)\geq(1+\phi)\rho\;.$
The term $(1-\rho)(1-\nu)$ can be viewed as the fraction of honest mining power discounted by the network delay $\Delta$. Therefore, this assumption requires that the honest mining power, even after accounting for network delay, exceeds the corrupt mining power by a small constant margin.
@@ A.4 Proof of Theorem 2
Given an execution of the Stubborn Nakamoto protocol, we can construct an execution of the Nakamoto [^75] We use the same confirmation depth $k$ as our Stubborn Nakamoto for the execution of the Nakamoto protocol. protocol. If the two executions are identical, in the sense that any honest node maintains the same finalized ledger at any point during the run, then Stubborn Nakamoto’s properties follow from those of Nakamoto. We show that under the network-aware honest majority assumption, the two executions are identical except for a negligible probability, and therefore the $T_{\rm conf}$-Liveness of Stubborn Nakamoto follows from the $T_{\rm conf}$-Liveness and consistency of Nakamoto (Pass et al.,, 2017; Shi,, 2022).
The reason the executions may differ is that an honest node $i$ may halt under Stubborn Nakamoto. Consider the first round $t_{1}$ such that some honest node $i$ certifies a block ${\sf B}$ and by round $t_{1}+4\Delta$ node $i$ halts having seen a $4\Delta$-potentially certified block ${\sf B}^{\prime}$ that conflicts with ${\sf B}$. (That is, node $i$ halted at either steps 1, 2 or 3)
Let $\sf Chain(\bar{{\sf B}})$ be the chain in node $i$’s local view in round $t_{1}$ containing ${\sf B}$ at least $k$-deep. Let ${\sf B}^{\prime}\in\mathcal{PC}_{i}^{t_{1}+4\Delta}(4\Delta)$ be a block that conflicts with ${\sf B}$, and let $t_{1}\leq t\leq t_{1}+4\Delta$ be the fist time at which ${\sf B}^{\prime}\in\mathcal{PC}_{i}^{t}(4\Delta)$. Let $\sf Chain(\bar{{\sf B}}^{\prime})$ be the chain in node $i$’s local record such that $h(\bar{{\sf B}}^{\prime})-h({\sf B}^{\prime})+1\geq k$ and $h(\bar{{\sf B}}^{\prime})\geq h(\tilde{{\sf B}})$ for any $\tilde{{\sf B}}\in{{\sf View}}_{i}^{t-4\Delta}$. That is, block ${\sf B}^{\prime}$ is at least $k$-deep in the chain $\sf Chain(\bar{{\sf B}}^{\prime})$, and the chain $\sf Chain(\bar{{\sf B}}^{\prime})$ is longer than any chain in the view of node $i$ as of $t-4\Delta$.
@@
Let ${\sf B}_{\rm common}$ the last block shared in common by $\sf Chain(\bar{{\sf B}})$ and $\sf Chain(\bar{{\sf B}}^{\prime})$. Suppose that ${\sf B}_{\rm common}$ is mined in round $s-1$. It holds that all blocks in $\sf Chain(\bar{{\sf B}})$ and $\sf Chain(\bar{{\sf B}}^{\prime})$ after ${\sf B}_{\rm common}$ are mined in round $s$ or later. Let $\tau=t-s$. Because the chains $\sf Chain(\bar{{\sf B}})$ and $\sf Chain(\bar{{\sf B}}^{\prime})$ differ by at least $k$ blocks, by total block upper bound (see Lemma A.3), it must be that $\tau>\frac{k}{2pn}$.
@@
Observe that it must be that ${\bf C}[s:t-5\Delta]\leq{\bf A}[s:t]$. Otherwise, there would have been a block ${\sf B}$ mined by an honest node during a convergence opportunity within $[s,t-4\Delta-\Delta]$, and that block ${\sf B}$ would have to appear in both $\sf Chain(\bar{{\sf B}})$ and $\sf Chain(\bar{{\sf B}})$ (see Section 17.3 of Shi, 2022). Therefore, it suffices to prove that, except with negligible $\lambda$ probability, it must be that ${\bf C}[s:t-5\Delta]>{\bf A}[s:t]$.
@@
${\bf C}[s:t-5\Delta]>(1-\varepsilon)(1-2pn\Delta)\alpha(\tau-5\Delta)\;,$
where $\alpha:=p\cdot(1-\rho)n$ denotes the expected number of honest nodes that mine a block in a given round. By adversarial block upper bound, (see Lemma A.2) we have that for any positive constant $\varepsilon_{a}$, it holds that
${\bf A}[s:t]<(1+\epsilon_{a})\beta\tau\;,$
@@
${\bf C}[s:t-5\Delta]$ $>(1-\varepsilon)(1-\nu)\alpha(\tau-5\Delta)$
@@
$>(1+\varepsilon^{\prime})\beta\tau$
@@
This implies that except with negligible in $\lambda$ probability we have that ${\bf C}[s:t-5\Delta]>{\bf A}[s:t]$, completing the proof.